Ceceila bought some 22 cent and some 17 cent stamps for the school office. Altogether she purchased 37 stamps for $7.54.How many stamps of each kind did she buy?
Answer by emargo19(101)

Let the quantity of 22 cents be x
Let the quantity of 17 cents be y
Therefore x+y=37 -----equation 1
Let the total value of 22 cents be 22x
And the total value of 17cents be 17y
Convert $7.54 into cents(multiply 7.54 *100=754)
Therefore 22x+17y=754------equation 2

x+y=37 ----------- equation 1
22x+17y=754 ------ equation 2
Use the Substitution method
Isolate variable x in equation 1
x=37-y
22(37-y)+17y=754
814-22y+17y=754
-22y+17y=754-814
-5y=-60
y=12
Substitute y = 12 into equation 1
x=37-12
x=25
therefore x=25
y=12
Ceceila bought 25, 22 cent and 12, 17 cent stamps for the school office